OpenStudy (anonymous):

2sqrt(11) times 8sqrt (22) @phi

OpenStudy (phi):

multiply them you can use the rule that \[ \sqrt{a}\cdot \sqrt{b}= \sqrt{a\cdot b} \]

OpenStudy (anonymous):

sqrt(242)

OpenStudy (phi):

don't forget the stuff out front also, it is best not to actually multiply out 11*22 because you should factor the 22 into 2*11, so you have 11*11*2 because you want to find pairs to take out of the square root.

OpenStudy (cggurumanjunath):